OpenStudy (anonymous1):

Could someone explain to me the following reasoning about derivatives? (more details in the reply) Could someone explain to me the following reasoning about derivatives? (more details in the reply) @Mathematics

OpenStudy (anonymous1):

It comes from the following expression: \[\frac{d}{dx}(\ln |u|)=\frac{1}{u}\frac{du}{dx}\] From the formula above we obtain the following formula by indefinite integration: \[\int\limits \frac{1}{u}du=\ln |u|+C\] (C is a constant) How is the result above obtained from the first expression?

OpenStudy (anonymous1):

I think I got it now, could someone confirm? \[d(\ln |u|)=\frac{1}{u}\frac{du}{dx}dx\] Using the following relationship: \[du=u'(x)dx=\frac{du}{dx}dx\] \[d(\ln |u|)=\frac{1}{u}du\] Integrating both sides: \[\int d(\ln |u|)=\int\frac{1}{u}du\] \[\ln |u|+C=\int\frac{1}{u}du\] which is the desired result. Is this correct?

myininaya (myininaya):

looks fine to me

OpenStudy (anonymous1):

Just one question: why are absolute value signs needed? Couldn't I write ln u instead of ln |u|?

OpenStudy (anonymous1):

In other words, could it be like below? \[\int \frac{1}{u}du=\ln u + C\]

OpenStudy (anonymous):

the absolutes are necessary because u must be positive - the log of a negative does not exist

OpenStudy (anonymous1):

But, if it is true that \[\frac{d(\ln u)}{du}=\frac{1}{u}\] then, why can't I make the reasoning below? \[d(\ln u) = \frac{1}{u}du\] \[\int d(\ln u) = \int\frac{1}{u}du\] \[\ln u+C = \int\frac{1}{u}du\] what is there in this manipulation that doesn't follow?

OpenStudy (zarkon):

let x>0 \[\frac{d}{dx}\ln(x)=\frac{1}{x}\] let x<0 \[\frac{d}{dx}\ln(-x)=\frac{1}{-x}(-1)=\frac{1}{x}\] thus \[\frac{d}{dx}\ln|x|=\frac{1}{x}\]

OpenStudy (anonymous1):

But what is wrong with the manipulation I did above?

OpenStudy (zarkon):

for what you have above that is only true if u>0

OpenStudy (zarkon):

unless you want your constant, C, to be complex

OpenStudy (anonymous1):

OK, I think I got it now.